# SoK: Large Language Model Copyright Auditing via Fingerprinting

### v2

@@ 6.2 Overview of LeaFBench
Based on the above summary, LeaFBench consists of three core components: a collection of base models and two categories of post-development techniques, which are integrated to enable diverse evaluation scenarios.
@@
Pre-trained Models. LeaFBench is built upon 7 mainstream pre-trained foundation models, including variants of Qwen [57], Llama [16, 58], Mistral [25], Gemma [56], and TinyLlama [80]. These models were selected for their popularity and advanced capabilities, with parameter counts ranging from 1.1B to 14B. Crucially, they all have a significant community presence and a large number of publicly available derivative models on platforms like Hugging Face, making them ideal for studying copyright lineage in the wild.
@@
Implementation of Parameter-altering Techniques. To simulate how adversaries adapt pre-trained base models, LeaFBench includes derivative models created using various widely-used parameter-altering techniques, including: (1) instruction tuning (IT), (2) general-purpose fine-tuning (FT), (3) parameter-efficient fine-tuning (PEFT), (4) quantization (QZ), (5) model merging (MM), and (6) distillation (DT). For each of the seven base models, we have collected no fewer than seven corresponding derivative models from Hugging Face that have been modified using one or more of these methods.
@@
Implementation of Parameter-independent Techniques. LeaFBench also evaluates the robustness of fingerprinting methods against common runtime parameter-independent techniques. This includes: (1) system prompts, covering general-purpose prompts (GP), role-playing prompts (RP), and chain-of-thought (CoT) prompts; (2) sampling strategies (SS), with configurations for greedy decoding, high-randomness, low-randomness, and standard settings; (3) LLM with RAG; and (4) adversarial manipulation (ADV), which considers two scenarios: rewriting the input query with a smaller LLM and perturbing the output logits.
@@
In total, LeaFBench comprises 149 distinct model instances. A comparison with existing studies is presented in Table 1. Arguably, given the complexity and rapid evolution of LLM post-development techniques, it is difficult for this benchmark to be fully exhaustive. Nonetheless, by focusing on a representative set of widely adopted techniques, LeaFBench strikes a balance between comprehensiveness and efficiency. This design enables it to serve as a practical tool for assessing the effectiveness and robustness of existing methods. Further details are provided in Appendix D.
@@
Table 1: Comparison of LeaFBench with the evaluation settings of representative existing studies.
@@
HuRef [76]
REEF [79]
LLMmap [44]
MET [13]
TRAP [17]
@@
Table 2: Overall performance comparison of different LLM fingerprinting methods.
@@
Table 3: Performance comparison of LLM fingerprinting methods across different source models. “PT” and “IT” refer to using the pre-trained models and instruction-tuned models as source models, respectively.
@@
6.3 Experimental Settings
Evaluated LLM Fingerprinting Methods. To provide a comprehensive evaluation, we select, design, and implement 8 representative fingerprinting methods that span the full taxonomy of our proposed framework. (1) For static fingerprinting, we evaluate two methods: HuRef [76] and Parameter Distribution Fingerprint (PDF) [75]. (2) For forward-pass fingerprinting, we include REEF [79]. (3) In the backward-pass fingerprinting, we implement a method named Gradient, which is inspired by the approach of TensorGuard [62] and incorporates several improvements. (4) For untargeted fingerprinting, we evaluate three black-box methods: LLMmap [44], Model Equality Testing (MET) [13], and our proposed baseline, $i.e.$, Sentence Embedding Fingerprinting (SEF). In general, SEF uses the averaged sentence embeddings of model responses as the fingerprint (see Appendix C for details). (5) Finally, for targeted fingerprinting, we select TRAP [17], which uses adversarial suffixes to generate unique query-response pairs. In particular, for all methods, we conduct at least 5 trials and report both the mean and the standard deviation.
Auditing Targets. We consider two types of source models for copyright auditing: pre-trained (PT) and instruction-tuned (IT) models. This choice reflects the most widespread and realistic auditing scenarios, where developers may wish to protect either their foundational pre-trained models or their more specialized, instruction-following variants.
  • Partial AUC with FPR[$\in[0,0.05]$ (pAUC): This metric evaluates performance in a low False Positive Rate range ($i.e.$, $[0,0.05]$). It highlights a method’s Uniqueness in avoiding false positives. This metric is standardized to $[0,1]$.
  • Mahalanobis Distance (MD): We use MD [63] to measure the Discriminability of a fingerprinting method by calculating the distance between the fingerprint distributions of derivative and independent models. A larger distance signifies a clearer separation between model lineages.
Table 4: Performance comparison of LLM fingerprinting methods across different parameter-altered techniques.
@@ 6.3 Experimental Settings
Table 5: Performance comparison of LLM fingerprinting methods across different parameter-independent techniques.
@@
6.4 Main Results
The results of our experiments are presented in Table 2 (Overall Performance), Table 3 (Performance Across Different Source Models), Table 4 (Robustness to Parameter-Altering Techniques), and Table 5 (Robustness to Parameter-Independent Techniques). In Table 2, we also report the overall accuracy (ACC) and the True Positive Rate at a 1% False Positive Rate (TPR@1%FPR), for reference. From these results, we distill several key takeaways regarding the effectiveness, robustness, and practical limitations of existing methods.
As shown in Table 2, white-box methods, which have direct access to model internals, achieve exceptional performance. Static methods like HuRef and PDF, for instance, achieve near-perfect AUC scores (0.994 and 0.995, respectively) and high MDs, indicating robust and highly discriminative fingerprints. In contrast, all evaluated black-box methods, although more practical, struggle to deliver reliable results, with AUC scores hovering in a much lower range (<<0.72). This performance gap stems from the nature of the features used: white-box methods can analyze a model’s parameters, whereas black-box methods rely on secondary, distinctive behaviors that are inherently more stochastic and less stable.
While white-box methods are broadly effective, as shown in Table 2, a clear performance hierarchy emerges. Specifically, static methods that directly analyze model weights ($i.e.$, HuRef, PDF) consistently deliver the highest accuracy and discriminability. Forward-pass and backward-pass methods, though still effective, show notably lower performance (AUCs of 0.896 and 0.798, respectively). This discrepancy likely arises from the immense scale of LLM parameter spaces, where independently trained models occupy highly distinct regions, making static weights inherently unique identifiers. In contrast, forward- and backward-pass methods rely on responses to a common input set, which can act as an alignment force. This shared context may induce more similar activation patterns or gradient landscapes across different models, thereby potentially reducing the fingerprint’s uniqueness compared to a direct analysis of the static parameters.
A compelling example is the comparison between the black-box methods TRAP and MET in Table 2. While TRAP shows a higher overall AUC (0.712 vs. 0.654), MET is arguably more practical, achieving a higher pAUC (0.672 vs. 0.625) for uniqueness and a higher MD (1.681 vs. 1.382) for discriminability. This demonstrates that reliance on a single headline metric can obscure critical operational characteristics. The prevailing emphasis on AUC and ACC in existing studies risks overestimating a method’s real-world applicability. For practical auditing, where avoiding false accusations (uniqueness) and ensuring a clear signal (discriminability) are vital, a more comprehensive evaluation is essential.
Our results in Table 3 reveal a key limitation of current black-box methods: their performance degrades substantially when the source model is a base pre-trained (PT) model rather than an instruction-tuned (IT) one. This discrepancy arises most probably because IT models, trained to follow instructions, exhibit more predictable and consistent response patterns, producing stronger and clearer fingerprint signals. In contrast, PT models generate more random and diverse outputs, resulting in weaker behavioral fingerprints that current methods struggle to detect reliably.
As shown in Table 4, white-box methods maintain high efficacy against most parameter-altering techniques, underscoring their general robustness. Static methods are particularly resilient across the board. However, their robustness is not absolute. For example, REEF and Gradient exhibit noticeable drops in AUC under fine-tuning and quantization. This indicates that while the model’s core identity is largely preserved, these modifications are still capable of altering the parameter distributions and representation geometry sufficiently to weaken the fingerprint’s signal strength.
Takeaway 6: Despite their broader applicability, black-box fingerprinting methods exhibit a critical lack of robustness to both parameter-altering and parameter-independent techniques, with the former posing a greater challenge.
The fragility of black-box fingerprinting is one of our key findings. Their dependence on secondary behaviors makes them highly susceptible to techniques that alter model outputs. Parameter-altering techniques are particularly damaging, often causing fingerprinting effectiveness to collapse, as shown in Table 4. In Table 5, parameter-independent techniques also degrade performance by manipulating the generation context, though typically less severely than direct weight modifications. This dual vulnerability constitutes the greatest barrier to the reliable deployment of black-box fingerprinting in real-world auditing scenarios.
@@ 6.5 Efficiency Evaluation
We evaluated the efficiency of 8 methods by measuring their total runtime in seconds. The detailed setting is in Appendix B. The results in Figure 4 reveal significant differences in computational cost. White-box methods are generally very efficient. Static and forward-pass approaches like PDF and REEF are the fastest, operating in under 50 seconds. The backward-pass method is moderately slower due to the added cost of backpropagation. In contrast, the efficiency of black-box methods varies dramatically. While standard untargeted methods like LLMmap and MET are reasonably fast, the targeted method TRAP is extremely time-consuming (nearly 11,000 seconds). This is due to its intensive optimization phase required to generate fingerprints. This indicates that the prohibitive computational cost of advanced black-box techniques may be a major barrier to their real-world application.
@@ 7 Insights, Challenges, and Future Directions
7.1 Toward Future Fingerprinting Techniques
Our analysis reveals that while white-box methods are highly effective, black-box methods still face significant challenges in reliability and robustness. To address these issues, we propose four promising future directions aimed at advancing black-box fingerprinting techniques:
  • Approximating White-Box Features. A promising direction is to develop black-box techniques that approximate or infer the powerful features used in white-box methods. For instance, future work could explore methods to reconstruct a model’s internal representation or distributions from its input-output behaviors to bridge the effectiveness gap.
  • Adopting Dynamic and Conversational Querying. Instead of relying on static, predefined query sets, future methods could employ dynamic, multi-turn conversational strategies. By adapting subsequent queries based on the model’s previous responses, an auditor could probe deeper and more distinctive characteristics.
  • Developing Hybrid Methods. Different black-box methods capture distinct aspects of a model’s behavior. A hybrid approach that combines multiple signals into a single, comprehensive fingerprint could offer greater resilience.
  • Balancing the Effectiveness-Efficiency in Targeted Methods. Targeted methods currently present a trade-off between effectiveness and computational cost, with high-AUC techniques like TRAP being impractically slow. Future research could focus on developing more efficient optimization algorithms that reduce the intensive computational overhead.
@@ 7.2 Beyond Existing Paradigms and Scope
Multi-model Fingerprinting. Existing LLM fingerprinting methods often target a single LLM. However, the landscape of modern AI applications is rapidly evolving toward complex, multi-model systems ($e.g.$, Mixture-of-Experts [5] and multi-agent system [18]) where several LLMs cooperate to complete tasks. This shift presents a fundamental challenge to current fingerprinting approaches. In a multi-model system, the final output is often a blending contribution of multiple distinct LLMs. This raises several open problems:
@@
  • Dynamic Interactions: In multi-model systems, the contribution of any single LLM can be dynamic and context-dependent. A fingerprinting technique may need to account for not just the static model but also its interactive behavior.
Therefore, future research may develop composition-aware fingerprinting techniques capable of isolating the statistical signatures of different models from a composite output. The goal is to design fingerprints robust not only to model modifications but also to the act of composition itself, ensuring reliable auditing in these collaborative AI systems.
Side-Channel Fingerprinting. Beyond analyzing a model’s direct outputs or internal parameters, an emerging and promising area of research is Side-Channel Fingerprinting. This approach extracts a fingerprint from an LLM’s runtime operational characteristics instead of its generated content. The core idea is that different LLM architectures exhibit unique, measurable side-channel signals during inference. For instance, recent studies have extracted fingerprints from memory usage patterns on edge devices [43] and the inter-token times of token generation in encrypted network traffic [1]. These signals could serve as a distinctive fingerprint for LLMs, even without access to the model’s parameters or its output.
A key advantage of this approach is its inherent robustness to content-altering modifications like paraphrasing, as it does not analyze the text itself. However, the field is still emerging. Future work may explore a wider range of side-channel signals ($e.g.$, GPU utilization, cache access patterns) and their robustness across diverse hardware and network conditions.
Audit Beyond Model Lineage. While much of the current research focuses on verifying a model’s origin or lineage, the principles of auditing can be extended to address a broader set of accountability challenges in LLM services. As LLMs are increasingly deployed in complex, opaque commercial services where users are billed for internal operations they cannot see, a new frontier for auditing emerges: verifying the fairness and honesty of the service itself. This is particularly relevant in scenarios where billing is based on complex, unobservable processes like multi-step reasoning, which creates incentives for dishonest practices. For example, a provider might maliciously inflate the number of “invisible” tokens to increase the final bill, or silently substitute a cheaper, lower-quality model while charging premium rates [52].

### v3

@@ 6.2 Construction of LeaFBench
Based on the above summary, LeaFBench consists of three core components: a collection of pre-trained models and two categories of post-development techniques, which are integrated to enable diverse evaluation scenarios.
@@
Pre-trained Models. LeaFBench is built upon 7 mainstream pre-trained foundation models, including variants of Qwen [qwen2.5], Llama [grattafiori2024llama3, touvron2023llama2], Mistral [mistral], Gemma [gemma2], and TinyLlama [zhang2024tinyllama]. These models were selected for their popularity and advanced capabilities, with parameter counts ranging from 1.1B to 14B. Crucially, they all have a widespread community with a large number of publicly available derivative models on platforms like Hugging Face, making them ideal for studying copyright lineage in the wild.
@@
Implementation of Parameter-altering Techniques. To simulate how adversaries adapt pre-trained models, LeaFBench includes derivative models created using various widely-used parameter-altering techniques, including: (1) instruction tuning (IT), (2) general-purpose fine-tuning (FT), (3) parameter-efficient fine-tuning (PEFT), (4) quantization (QZ), (5) model merging (MM), and (6) distillation (DT). For each of the seven pre-trained models, we have collected no fewer than seven corresponding derivative models from Hugging Face that have been modified using one or more of these methods.
@@
Implementation of Parameter-independent Techniques. LeaFBench also evaluates the robustness of fingerprinting methods against common runtime parameter-independent techniques. This includes: (1) system prompts, covering general-purpose prompts (GP), role-playing prompts (RP), and chain-of-thought (CoT) prompts; (2) sampling strategies (SS), with configurations for greedy decoding, high-randomness, low-randomness, and moderate settings; (3) LLM with RAG; and (4) adversarial manipulation (ADV), which considers two scenarios: rewriting the input query with a smaller LLM and perturbing the output logits with random noise.
Organizing the Model Instances into Triplets. In total, LeaFBench includes 149 distinct model instances. To operationalize our evaluation, we structure LeaFBench to directly address the copyright auditing problem formalized in Definition 1. Specifically, we construct a comprehensive dataset in which each sample is represented as a triplet $(M_{o}^{i},M_{s}^{i},y^{i})$. Here, $M_{o}^{i}$ denotes the source model, selected from a pool of 13 base models, including both pre-trained (PT) and instruction-tuned (IT) variants. This design captures realistic auditing scenarios, where developers may seek to protect either their foundational pre-trained models or their specialized instruction-following derivatives. Besides, the suspicious model $M_{s}^{i}$ is drawn from the complete set of 149 model instances, encompassing all base models and their derivatives. The label $y^{i}\in\{0,1\}$ serves as the ground truth for the auditing task, where $y^{i}=1$ if and only if $M_{s}^{i}$ is a known derivative of $M_{o}^{i}$, and $y^{i}=0$ otherwise. This formulation results in a total of $13\times 149=1937$ evaluation samples within the designed LeaFBench.
@@
A comparison between existing studies and LeaFBench is presented in Table I. Notably, given the complexity and rapid evolution of LLM post-development techniques, it is difficult for this benchmark to be fully exhaustive. Nonetheless, by focusing on a representative set of widely adopted techniques, LeaFBench strikes a balance between comprehensiveness and efficiency. This design enables it to serve as a practical tool for assessing the effectiveness and robustness of existing methods. Further details are provided in Appendix .4.
@@
TABLE I: Comparison of LeaFBench with the evaluation settings of representative existing studies. indicates the technique is included, while indicates it is not.
@@
HuRef [zeng2024huref]
REEF [zhang2025reef]
PDF [yoon2025intrinsic]
LLMmap [pasquini2025llmmap]
MET [gao2025model]
TRAP [gubri2024trap]
@@
TABLE II: Overall performance comparison of different LLM fingerprinting methods.
@@
6.3 Experimental Settings
Evaluated LLM Fingerprinting Methods. To provide a comprehensive evaluation, we select, design, and implement 8 representative fingerprinting methods that span the full taxonomy of our proposed framework. (1) For static fingerprinting, we evaluate two methods: HuRef [zeng2024huref] and Parameter Distribution Fingerprint (PDF) [yoon2025intrinsic]. (2) For forward-pass fingerprinting, we include REEF [zhang2025reef]. (3) In the backward-pass fingerprinting, we implement a method named Gradient, which is inspired by the approach of TensorGuard [wu2025gradient] and incorporates several improvements. (4) For untargeted fingerprinting, we evaluate three black-box methods: LLMmap [pasquini2025llmmap], Model Equality Testing (MET) [gao2025model], and our proposed simple baseline, Sentence Embedding Fingerprinting (SEF), inspired by [sun2025idiosyncrasies]. In general, SEF uses the averaged sentence embeddings of model responses as the fingerprint (see Appendix .3 for details). (5) Finally, for targeted fingerprinting, we select TRAP [gubri2024trap], which uses adversarial suffixes to generate unique query-response pairs. In particular, for all methods, we conduct at least 5 trials and report both the mean and the standard deviation.
  • Partial AUC with FPR[$\in[0,0.05]$ (pAUC): This metric evaluates performance in a low False Positive Rate (FPR) range ($i.e.$, $[0,0.05]$). It highlights a method’s Uniqueness in avoiding false positives. Notably, this metric is standardized to $[0,1]$ in our experiments.
  • Mahalanobis Distance (MD): MD [xiang2008learning] is to measure the Discriminability of a fingerprinting method by calculating the distance between the fingerprint distributions of derivative and independent models. A larger distance signifies a clearer separation between model lineages.
TABLE III: Performance comparison of LLM fingerprinting methods across different source models. “PT” and “IT” refer to using the pre-trained models and instruction-tuned models as source models, respectively.
@@ 6.3 Experimental Settings
TABLE IV: Performance comparison of LLM fingerprinting methods across different parameter-altered techniques.
@@
6.4 Main Results
The results of our experiments are presented in Table II (Overall Performance), Table III (Performance Across Different Source Models), Table IV (Robustness to Parameter-Altering Techniques), and Table V (Robustness to Parameter-Independent Techniques). In Table II, we also report the overall accuracy (ACC) and the True Positive Rate at a 1% False Positive Rate (TPR@1%FPR), for reference. From these results, we distill several key takeaways regarding the effectiveness, robustness, and practical limitations of existing methods.
As shown in Table II, white-box methods, which have direct access to model internals, achieve exceptional performance. Static methods like HuRef and PDF, for instance, achieve near-perfect AUC scores (0.994 and 0.995, respectively) and high MDs, indicating robust and highly discriminative fingerprints. In contrast, all evaluated black-box methods, although more practical, struggle to deliver reliable results, with AUC scores hovering in a much lower range (<<0.72). This performance gap stems from the nature of the features used: white-box methods can analyze a model’s parameters, whereas black-box methods rely on secondary behaviors that are inherently more stochastic and less stable.
While white-box methods are broadly effective, as shown in Table II, a clear performance gap emerges. Specifically, static methods that directly analyze model weights ($i.e.$, HuRef, PDF) consistently deliver the highest accuracy and discriminability. Forward-pass and backward-pass methods, though still effective, show lower performance (AUCs of 0.896 and 0.798, respectively). This discrepancy likely arises from the immense scale of LLM parameter spaces, where independently trained models have highly distinct parameters, making static weights inherently unique identifiers. In contrast, forward-pass and backward-pass methods rely on activations and responses to a common input set. This shared context may induce more similar activation patterns or gradient landscapes across different models, thereby potentially reducing the fingerprint’s uniqueness compared to a direct analysis of the static parameters.
A compelling example is the comparison between the black-box fingerprinting methods TRAP and MET in Table II. While TRAP shows a higher overall AUC (0.712 vs. 0.654), MET is arguably more practical, achieving a higher pAUC (0.672 vs. 0.625) for uniqueness and a higher MD (1.681 vs. 1.382) for discriminability. This demonstrates that reliance on a single headline metric can obscure critical operational characteristics. The prevailing emphasis on AUC and ACC in existing studies risks overestimating a method’s real-world applicability. For practical auditing, where avoiding false accusations (uniqueness) and ensuring a clear signal (discriminability) are vital, a more comprehensive evaluation is essential.
Our results in Table III reveal a key limitation of current black-box methods: their performance degrades substantially when the source model is a base pre-trained (PT) model rather than an instruction-tuned (IT) one. This discrepancy arises most probably because IT models, trained to follow instructions, exhibit more predictable and consistent response patterns, producing stronger and clearer fingerprint signals. In contrast, PT models generate more random and diverse outputs, resulting in weaker behavioral fingerprints that current methods struggle to detect reliably.
As shown in Table IV, white-box methods maintain high efficacy against most parameter-altering techniques, underscoring their general robustness. Static methods are particularly resilient across the board. However, their robustness is not absolute. For example, REEF and Gradient exhibit noticeable drops in AUC under fine-tuning and quantization. This indicates that while the model’s core identity is largely preserved, these modifications are still capable of altering the parameter distributions and representation geometry sufficiently to weaken the fingerprint’s signal strength.
Takeaway 6: Despite their broader applicability, black-box methods exhibit a critical lack of robustness to both parameter-altering and parameter-independent techniques, with the former posing a greater challenge.
The fragility of black-box fingerprinting is one of our key findings. Their dependence on secondary behaviors makes them highly susceptible to techniques that alter model outputs. Parameter-altering techniques are particularly damaging, often causing fingerprinting effectiveness to collapse, as shown in Table IV. In Table V, parameter-independent techniques also degrade performance by manipulating the generation context, though typically less severely than direct weight modifications. This dual vulnerability constitutes the greatest barrier to the reliable deployment of black-box fingerprinting in real-world auditing scenarios.
TABLE V: Performance comparison of LLM fingerprinting methods across different parameter-independent techniques.
@@ 6.5 Efficiency Evaluation
In this section, we evaluated the efficiency of 8 fingerprinting methods by measuring their total runtime in seconds. To streamline the efficiency evaluation process while ensuring a fair and consistent comparison, we conducted the runtime tests on a curated subset of four models from the Qwen2.5-7B lineage. All experiments were uniformly executed on a standardized hardware setup consisting of two NVIDIA RTX 4090 GPUs. All other hyperparameters and settings for the evaluated fingerprinting methods were kept at their default values to maintain consistency with their original implementations.
The results in Figure 4 reveal significant differences in computational cost. White-box methods are generally very efficient. Static and forward-pass approaches like PDF and REEF are the fastest, operating in under 50 seconds. The backward-pass method is moderately slower due to the added cost of backpropagation. In contrast, the efficiency of black-box methods varies dramatically. While standard untargeted methods like LLMmap and MET are reasonably fast, the targeted method TRAP is extremely time-consuming (nearly 11,000 seconds). This is due to its intensive optimization phase required to generate fingerprints. This indicates that the prohibitive computational cost of advanced black-box techniques may be a major barrier to their real-world application.
@@ 7 Insights, Challenges, and Future Directions
7.1 Towards Future Fingerprinting Techniques
Our analysis reveals that while white-box methods are highly effective, black-box methods remain unreliable and fragile (Takeaway 1&6). To address these issues, we propose four promising future directions aimed at advancing black-box fingerprinting techniques:
  • Approximating White-Box Features. Our study confirms that static parameters provide the most robust fingerprints (Takeaway 2). A critical future direction is to develop black-box techniques that can infer or reconstruct these white-box features ($e.g.$, internal representation geometry or parameter statistics) from observable input-output behaviors, thereby bridging the effectiveness gap.
  • Adopting Dynamic and Conversational Querying. We found that black-box methods fail when auditing PT models, which produce more diverse and random outputs than IT models (Takeaway 4). This necessitates a shift from static query sets to dynamic and conversational querying strategies capable of eliciting stable, core characteristics from these less predictable base models.
  • Developing Hybrid Methods. Different black-box methods capture distinct aspects of a model’s behavior (Takeaway 3&6). A hybrid approach that combines multiple signals into a single, comprehensive fingerprint could offer greater resilience.
  • Balancing the Trade-off between Effectiveness and Efficiency. We observed a critical trade-off where methods are either impractically slow ($e.g.$, TRAP) or, if efficient, show poor overall performance despite better reliability metrics ($e.g.$, MET) (Takeaway 3). Future research could explore co-optimizing for effectiveness, efficiency, and practical reliability ($i.e.$, low false-positives and high discriminability), not just headline accuracy.
@@ 7.2 Beyond Existing Paradigms and Scope
Multi-model Fingerprinting. Existing LLM fingerprinting methods often target a single LLM. However, the landscape of modern AI applications is rapidly evolving towards complex, multi-model systems ($e.g.$, Mixture-of-Experts [cai2025survey] and multi-agent system [he2025routemark]) where several LLMs cooperate to complete tasks. In a multi-model system, the final output is often a blending contribution of multiple distinct LLMs. This shift raises several fundamental challenges to LLM fingerprinting methods:
@@
  • Dynamic Interactions: In multi-model systems, the contribution of any single LLM can be dynamic and context-dependent. A multi-model fingerprinting technique may need to account for not just a static model but also its interactive behavior.
Therefore, future research may develop composition-aware fingerprinting techniques capable of isolating the statistical signatures of different models from a composite output. The goal is to design fingerprints robust not only to model modifications but also to the act of composition, ensuring reliable auditing in these collaborative AI systems.
Side-Channel Fingerprinting. Beyond analyzing a model’s direct outputs or internal parameters, an emerging and promising area of research is Side-Channel Fingerprinting. This approach extracts a fingerprint from an LLM’s runtime operational characteristics instead of its generated content. The core idea is that different LLM architectures exhibit unique, measurable side-channel signals during inference. For instance, recent studies have extracted fingerprints from memory usage patterns on edge devices [nazari2024llm] and the inter-token times of token generation in encrypted network traffic [alhazbi2025llms]. These signals could serve as a distinctive fingerprint for LLMs, even without access to the model’s parameters or its output.
A crucial advantage of side-channel fingerprinting is its inherent robustness to content-altering modifications like paraphrasing, as it does not analyze the text itself. However, the field is still emerging. Future work may explore a wider range of side-channel signals ($e.g.$, GPU utilization, cache access patterns) and their robustness across diverse hardware and network conditions.
Audit Beyond Model Origin. While much of the current research focuses on verifying a model’s origin or lineage, the principles of auditing can be extended to address a broader set of accountability challenges in LLM services. As LLMs are increasingly deployed in complex, opaque commercial services where users are billed for internal operations they cannot see, a new frontier for auditing emerges: verifying the fairness and honesty of the service itself. This is particularly relevant in scenarios where billing is based on complex, unobservable processes like multi-step reasoning, which creates incentives for dishonest practices. For example, a malicious provider might maliciously inflate the number of “invisible” tokens to increase the final bill [sun2025invisible], or silently substitute a cheaper, lower-quality model while charging premium rates [gao2025model, cai2025you].
@@
7.3 Evasion Attacks against LLM Fingerprinting
While our experiments in LeaFBench evaluate robustness against common model modifications and two straightforward adversarial manipulations, a malicious adversary may actively attempt to evade detection using evasion attacks. Recent studies have begun to explore this adversarial dimension. Nasery et al. [nasery2025robust] present four fundamental vulnerabilities in existing black-box fingerprinting methods: verbatim verification, overconfidence, unnatural queries, and statistical signatures. They further demonstrate that adaptive attacks could bypass existing auditing with minimal impact on model utility. Other works propose specific evasion techniques. For example, [zeng2024huref] notes that white-box fingerprints could be vulnerable to weight rearrangement attacks like permutation, and [kurian2025attacks] shows that using an auxiliary LLM to paraphrase the output can effectively obfuscate a model’s identity against black-box fingerprinting.
Despite these initial efforts, research into evasion attacks remains underexplored. Most existing attacks target specific vulnerabilities in either white-box or black-box methods, lacking a comprehensive understanding. Existing attacks also could not achieve both a low latency and a high attack effectiveness [nasery2025robust, zeng2024huref, shao2025reading]. Future research could explore more sophisticated and generalizable evasion strategies, such as combining multiple evasion techniques. Another promising direction is to develop “removal” techniques that selectively erase fingerprint signals from a model’s parameters with minimal degradation to its overall performance, posing a significant threat to all fingerprinting paradigms. Studying more advanced evasion attacks may lead to a deeper understanding of these adversarial threats, which is crucial for developing truly robust fingerprinting solutions.